OpenStudy (anonymous):

Find the exact positive value of c if the area between the graph of y = x^2 − c^2 and the x axis is 40.

OpenStudy (anonymous):

any help would be appreciated!! :/

myininaya (myininaya):

i don't get what the limits are

OpenStudy (anonymous):

same. I'm so confused

myininaya (myininaya):

oh i guess we are using the intersections...

myininaya (myininaya):

that is the intersections of the y=x^2-c^2 and the x-axis

myininaya (myininaya):

x^2-c^2=0 when x=?

OpenStudy (anonymous):

sqrt (c^2)

myininaya (myininaya):

so we have our upper and lower limit there x^2=c^2 x=sqrt(c^2) x=c or -c c is the upper limit -c is the lower limit \[\int\limits_{-c}^{c}(0-(x^2-c^2))=40\]

myininaya (myininaya):

solve that for c

OpenStudy (anonymous):

so solve the integral for c?

myininaya (myininaya):

 c^2 there is the y-intercept of y=x^2-c^2